# PASS-FC: Progressive and Adaptive Search Scheme for Fact Checking of Comprehensive Claims

### v1

@@ A.3 Label Standardization
A.4 Claim Verification Results on Factbench
FacTool-QA FELM-WK Factcheck-GPT Average
Framework Base Model F1 Acc F1 Acc F1 Acc F1 Acc
Perplexity.ai Sonar-online 0.640 0.815 0.570 0.689 0.680 0.705 0.630 0.736
FactScore GPT-3.5-Turbo 0.540 0.584 0.565 0.644 0.585 0.631 0.563 0.620
FacTool GPT-3.5-Turbo 0.625 0.650 0.560 0.606 0.635 0.686 0.607 0.647
FacTool (Rerun) GPT-3.5-Turbo 0.590 0.605 - - - - - -
Factcheck-GPT GPT-4-Turbo 0.755 0.818 0.665 0.763 0.710 0.674 0.710 0.751
Factcheck-GPT (Rerun) GPT-4-Turbo 0.696 0.751 - - - - - -
SAFE GPT-3.5-Turbo 0.591 0.627 0.465 0.466 0.572 0.550 0.542 0.547
PASS-FC GPT-3.5-Turbo 0.672 0.785 0.586 0.704 0.698 0.757 0.652 0.749
PASS-FC GPT-4omini 0.770 0.811 0.659 0.694 0.701 0.720 0.710 0.742
PASS-FC GPT-4o 0.828 0.863 0.681 0.708 0.737 0.743 0.749 0.771
PASS-FC GPT-4-Turbo 0.810 0.854 0.698 0.752 0.734 0.761 0.748 0.789
Table 4: Macro-F1 and accuracy for the fact-checking task on the Factbench. The highest results are in bold, and the second-best results are underlined. Results within the shaded area are cited from OpenFactCheck ((Wang et al., 2024b)).
Table 4 demonstrate that PASS-FC consistently outperforms other frameworks when utilizing the same large language model. This superiority is particularly evident when comparing PASS-FC with GPT-4-Turbo to Factcheck-GPT with GPT-4-Turbo, which represents the best-performing baseline.
PASS-FC with GPT-4-Turbo achieves higher scores across all datasets and metrics compared to Factcheck-GPT with GPT-4-Turbo. This performance advantage is observed in the FacTool-QA, FELM-WK, and Factcheck-GPT datasets, as well as in the overall average scores.
The replication attempts for FacTool and Factcheck-GPT failed to reproduce the performance reported in the original studies. This discrepancy can be attributed to two main factors:
Firstly, the inherent variability and temporal evolution of search engine rankings contribute to inconsistencies in information retrieval. This variability introduces an element of randomness in the fact-checking process, potentially affecting the models’ performance.
Secondly, and more significantly, these models do not account for the temporal context of claims, specifically the verification time metadata. The FacTool-QA dataset, for instance, lacks this crucial temporal information. PASS-FC addresses this limitation by inferring a verification time based on the dataset’s publication date on arXiv (2023-07-26), assuming all claims were verified prior to this date.
This temporal consideration is critical because the veracity of claims can change over time. A pertinent example from the FacTool-QA dataset illustrates this point: the claim "The United States has 94 operating reactors" may have been accurate at the time of dataset annotation in 2023. However, recent developments, such as the commencement of commercial operations at Vogtle Unit 4 in Georgia on April 29, 2024, have altered this fact [^6] https://www.eia.gov/tools/faqs/faq.php?id=228&t=21.
Consequently, models evaluating this claim in 2024 without considering the original verification time might produce results that are factually correct for the current time but inconsistent with the dataset’s ground truth labels. This temporal discrepancy underscores the importance of incorporating time-aware fact-checking mechanisms.
A.5 Pseudocode for PASS-FC Pipeline
Algorithm 1 PASS-FC Pipeline Algorithm
1:Query $q$, Response $r$, Metadata $m$, Source Language $l$
2:Veracity Label $v$
3:$claims\leftarrow\texttt{ClaimDetection}(q,r,m)$
4:for each $claim\in claims$ do
5:   $c\leftarrow\texttt{ClaimAugmentation}(claim)$
6:   $iteration\leftarrow 0$
7:   $evidence\leftarrow\emptyset$
8:   $history\leftarrow\emptyset$Initialize verification history}
9:   while $iteration<\mathrm{MAX\_ITERATIONS}$ do
10:      if $\mathrm{need\_new\_queries}$ then
11:         $queries\leftarrow\texttt{QueryGeneration}(c,l,history)$History-aware query generation}
12:         $evidence\leftarrow\texttt{EvidenceRetrieval}(queries)$
13:      end if
14:      $v\leftarrow\texttt{ClaimVerification}(c,evidence,history)$History-informed verification}
15:      $(feedback,continue,need\_new\_queries)\leftarrow\texttt{Reflection}(c,evidence,v,history)$
16:      $history\leftarrow history\cup\{(c,v,feedback,iteration)\}$Append complete context}
17:      if $\neg continue$ then
18:         break
19:      end if
20:      $iteration\leftarrow iteration+1$
21:   end while
22:end for
23:return $\texttt{AggregateLabels}(history)$Final decision from history}
A.6 Example Prompts
@@
[Examples]
<context>: Who won the match between Tomas Berdych and Gael Monfis on Monte Carlo Masters, 2015?
<passage>: Tomas Berdych defeated Gael Monfis 6-1, 6-4 on Saturday. The sixth-seed reaches Monte Carlo Masters final for the first time. Berdych will face either Rafael Nadal or Novak Djokovic in the final.
<response>: [{{"claim": "Tomas Berdych defeated Gael Monfis 6-1, 6-4"}}, {{"claim": "Tomas Berdych defeated Gael Monfis 6-1, 6-4 on Saturday"}}, {{"claim": "Tomas Berdych reaches Monte Carlo Masters final"}}, {{"claim": "Tomas Berdych is the sixth-seed"}}, {{"claim": "Tomas Berdych reaches Monte Carlo Masters final for the first time"}}, {{"claim": "Berdych will face either Rafael Nadal or Novak Djokovic"}}, {{"claim": "Berdych will face either Rafael Nadal or Novak Djokovic in the final"}}]
<context>: How many photos does Tinder show by default, and can users access additional photos beyond this limit?
<passage>: Tinder only displays the last 34 photos - but users can easily see more. Firm also said it had improved its mutual friends feature.
<response>: [{{"claim": "Tinder only displays the last photos"}}, {{"claim": "Tinder only displays the last 34 photos"}}, {{"claim": "Tinder users can easily see more photos"}}, {{"claim": "Tinder said it had improved its feature"}}, {{"claim": "Tinder said it had improved its mutual friends feature"}}]
@@
[Definitions about Fact]
[Instructions]
1. You are given a passage, a factual claim extracted from the passage, and a text feedback about the problem of the extracted claim. If the claim can be completely decomposed into multiple atomic factual claims, your task is to break the given factual claim down into a list of atomic factual claims, based on the given [Definitions about Fact]. Otherwise, return "None".
3. If the feedback contains rational suggestions, they should be adopted to refine the final result.
4. Extract clear, unambiguous atomic factual claims to check from the given claim, avoiding vague references like he’, she’, it’, or this’, and using complete names.
[Input Format Instruction]
<context>: Context for <passage> to help you understand it better.
<passage>: The passage where the following claim is extracted from.
<extracted_claim>: A claim that was extracted from the passage in a former round of claim extraction.
<feedback>: Feedback about the former round of claim extraction. It may contain problems about the extracted claim and corresponding suggestions.
[Output Format Instruction]
1. If the <extracted_claim> is already an atomic factual claim which cannot be further decomposed, output "None".
2. Otherwise, your response MUST be a list of dictionaries. Each dictionary should contains the key "claim", which correspond to the extracted claim (with all coreferences resolved).
[response format]
[Examples]
{examples}
Now complete the following, ONLY RESPONSE IN A LIST OF DICTIONARIES FORMAT, NO OTHER WORDS!!!:
<context>: {prompt}
<passage>: {input}
<extracted_claim>: {claim}
<feedback>: {feedback}
<response>:
Figure 6: The prompt used in Reflected Claim Decomposition.
@@ 3. Requests for resolving vague references:
Figure 7: The prompt used in Claim Augmentation.
@@
Figure 8: The prompt used in Advanced Search.
[Instructions]
2. Only provide domain suffixes, not full URLs.
3. Include reliable sources such as:
Government and official websites (.gov, .org)
Encyclopedia websites (.wiki)
Reputable news outlets (provide their official domain names)
6. In summary, for each claim, provide:
Recommended domain suffixes for searching
Domains to exclude (marked with a minus sign)
[response format]: [’url1’, url2’, ’-url3’]
[Examples]
<CLAIM>: The Eiffel Tower was built in 1889.
<CLAIM>: COVID-19 vaccines are safe and effective.
<CLAIM>: {input}
<HISTORY>: {feedback}
<RESPONSE>:
Figure 9: The prompt used in Site-Restricted Search.
[Supported Languages]
"Dutch", "Estonian", "Filipino", "Finnish", "French (Canada)",
"Swedish", "Thai", "Turkish", "Ukrainian", "Vietnamese", "Zulu"]
[Instructions]
[response format]: [’language1’, language2’]
[Examples]
<RESPONSE>: ["German"]
<RESPONSE>: ["Korean"]
<RESPONSE>: ["Portuguese (Brazil)"]
<RESPONSE>: ["Russian", "Ukrainian"]
<RESPONSE>: None
Now complete the following:
<CLAIM>: {input}
<RESPONSE>:
Figure 10: The prompt used in Multilingual Search.
You are given a piece of factual claim. Your task is to identify whether there are any factual errors within the claim. Besides a claim, background information about the claim such as its valid time and descriptions of potentially confusing named entities is also provided. Note that the background information is only for better understanding of the claim. You SHOULD NOT judge the factuality of the background information.
Some evidence has already been retrieved in previous attempts to verify the claim, and several(at least one) verifications were made. The previous series of actions taken are saved in <STEP_HISTORY>. However, upon further reflection, it was discovered that there were issues with the previous verification process, and negative <FEEDBACK> was provided. The <FEEDBACK> includes an analysis of the problems with the previous attempt and a plan for the direction of future steps.
Based on the <FEEDBACK>, additional evidence may have been retrieved. If provided, please make good use of this <NEW_EVIDENCE>, along with the previous verification process stored in <STEP_HISTORY>, to verify the claim. If only the <FEEDBACK> is provided without new evidence, it means that the existing evidence in <STEP_HISTORY> is sufficient. You should follow the instruction in the <FEEDBACK> to identify and consider relevant evidence for re-verification.
When you are judging the factuality of the given text, you could reference the provided evidences if needed. The provided evidences may be helpful. You must be careful when using the evidence to judge the factuality of the given text. Supportive evidence should be consistent with the claim in terms of the facts, time, and named entities. Some evidence may contradict each other. But good evidence should comes from a credible source. If there is insufficient evidence to either support or refute the claim, classify it as INCONCLUSIVE.
The response should be a dictionary with four keys - "reasoning", "factuality", "error", and "correction", which correspond to the reasoning, whether the given text is factual or not (Boolean - True, False, or INCONCLUSIVE), the factual error present in the text, and the corrected text.
<CLAIM>:
{claim}
<STEP_HISTORY>:
{steps}
<FEEDBACK>:
{feedback}
<NEW_EVIDENCE>:
{evidence}
<OUTPUT>:
Figure 11: The prompt used in Claim Verification.

### v2

@@ A.4 Label Standardization
Frameworks Base Model Source F1 Accuracy
DEFAME GPT-4o-0806 AVeriTeC KB N/A 74.0
DEFAME GPT-4omini-2024-07-18 AVeriTeC KB N/A 59.2
Papelo (first 200) GPT-4o-0513 google serper 77.9 75.0
PASS-FC GPT-4o-0806 google serper 77.6 72.0
PASS-FC GPT-4omini-2024-07-18 google serper 69.9 70.2
PASS-FC GPT-35-turbo-0613 google serper 66.6 65.8
DEFAME (first 220) GPT-4o-0806 google serper 71.0 61.1
PASS-FC (first 220) GPT-4o-0806 google serper 72.6 67.3
Table 6: Performance metrics across different framework and model combinations on AVeriTeC-Dev. Results within the shaded area are cited from their original paper (Braun et al., 2024; Malon, 2024). Note that Papelo only report results on the first 200 sample of AVeriTeC-Dev and we put the best result amoung their 9 settings here (They treat AVeriTeC-Dev as the development set.)
A.5 AVeriTeC Challenge Frameworks Comparison
Most systems in the AVeriTeC Challenge used AVeriTeC’s official knowledge base, which includes all golden evidence plus 1000 irrelevant pieces per claim, simplifying the task. The Papelo work (Malon, 2024) is one of only three that used Google search directly (also the best performing one amoung the three), but it is explicitly stated by its authors [^6] https://github.com/cdmalon/fever2024 to be specifically designed for AVeriTeC, not for general fact-checking. This situation remains for other frameworks except DEFAME. Therefore, we did not include them in our main experiment comparisons as our focus is on general-purpose fact-checking systems.
We report comparisons with DEFAME and Papelo in Table 6. DEFAME’s performance varies significantly across base models, suggesting heavy reliance on model capabilities. We tested DEFAME with Google search on only the first 200 examples (cost around 5 hours and 50\$). Even with this limited sample, our system outperformed DEFAME using the same search source. Papelo works better on their reported result. But they treat AVeriTeC-Dev as the development set. We put the best result amoung their 9 settings here.
Framework no ru es it id pl pt ro sr tr Overall
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
FacTool 62.0 62.0 57.7 57.0 56.5 58.0 52.8 52.0 68.4 66.0 51.2 54.0 54.8 55.0 53.3 55.0 51.0 55.0 55.3 58.0 56.9 57.2
FIRE-4o 65.2 67.0 62.2 64.0 56.6 60.0 47.0 48.0 73.8 74.0 60.4 64.0 56.9 58.0 56.3 58.0 53.6 57.0 58.0 60.0 59.1 61.0
FIRE-4omini 64.8 67.0 60.1 63.0 49.5 54.0 51.3 52.0 73.3 73.0 60.4 64.0 51.9 52.0 62.8 63.0 50.8 53.0 59.1 63.0 58.9 60.4
passfc-35 81.9 82.0 66.6 67.0 56.7 59.0 53.1 54.0 73.0 73.0 53.7 55.0 58.0 58.0 54.9 55.0 53.6 57.0 60.8 61.0 61.6 62.1
passfc-4o 80.7 81.0 75.2 76.0 63.9 66.0 54.6 55.0 76.9 77.0 66.9 68.0 60.8 61.0 40.7 45.0 52.4 57.0 59.1 63.0 63.7 64.9
passfc-4omini 80.3 81.0 66.9 69.0 52.5 58.0 46.3 48.0 70.5 71.0 63.0 67.0 65.8 66.0 46.4 50.0 39.9 49.0 54.0 60.0 59.3 61.9
- XLE 77.1 78.0 58.2 62.0 51.8 58.0 49.5 51.0 70.3 71.0 60.7 65.0 63.3 64.0 46.5 51.0 40.9 49.0 47.6 56.0 57.4 60.5
- lang adap 78.5 79.0 67.6 69.0 57.4 61.0 52.4 53.0 69.8 70.0 64.0 67.0 56.0 57.0 47.9 52.0 46.2 54.0 59.9 64.0 60.6 62.6
Average 73.8 74.6 64.3 65.9 55.6 59.3 50.9 51.6 72.0 71.9 60.0 63.0 58.4 58.9 51.1 53.6 48.6 53.9 56.7 60.6 - -
Table 7: Framework performance across different languages on X-FACT. "XLE" and "lang adap" are short for "Cross-Lingual Expansion" and "language adaptation". The last row shows the average result of every column.
Framework Base Model Money (\$)
FacTool gpt-35-turbo-0613 4.31
SAFE gpt-35-turbo-0613 2.89
FactCheck-GPT gpt-4omini-0718 2.31
FIRE gpt-4omini-0718 0.27
FIRE gpt-4o-0806 4.55
PASS-FC gpt-4o-0806 4.74
PASS-FC gpt-4omini-0718 0.19
PASS-FC gpt-35-turbo-0613 0.58
Table 8: Total API costs in USD for LLM token consumption across different framework and model combinations. The monetary values represent the total cost of API calls measured during experimental runs on AVeriTeC-Dev and ScirFact-Open.
A.6 Case Study
We provide two compact examples—one for Temporal Grounding (Figure 6) and one for XLE (Figure 7)—to illustrate how PASS-FC works in practice. In the second example, XLE generates Spanish and Portuguese queries that surface new evidence, altering the final verdict. To save space, we keep only the essential steps (e.g., the evidence passages actually used in validation). Full, unabridged traces are available in the supplementary material along with all experimental results.
"step_num": 1,
"action": "contextual_grounding",
"original_claim": "The United States has 94 operating reactors"
[…]
"step_num": 2,
"action": "generate_query",
"result": {
"advanced query": [
"United States nuclear reactors count 2023", "How many operating nuclear reactors are in the United States?"]}
"step_num": 3
"action": "EvidenceSearch",
"result": {
"United States nuclear reactors count 2023": [
[…]
"How many operating nuclear reactors are in the United States?": [
[…]
"step_num": 4
"action": "factuality_check",
"reasoning": "The claim states that the United States has 94 operating reactors. However, multiple credible sources indicate that as of August 1, 2023, the number of operating commercial nuclear reactors is 93. For instance, the U.S. Energy Information Administration (EIA) confirms that there are 93 operating reactors. Additionally, Statista also reports 93 operable units for 2023. While some sources mention 94 reactors, they may be referencing a future projection or a different context, which does not align with the current operational status as of the specified date.",
"error": "The claim incorrectly states the number of operating reactors as 94 instead of the accurate figure of 93.",
"correction": "The United States has 93 operating reactors.",
"factuality": false
Figure 6: A simplified case due to space restriction. It shows how temporal grounding works by discovering the claim period. Complete cases can be found in our supplementary materials, where we append all of the experiment results.
"step_num": 1,
"action": "contextual_grounding",
"original_claim": "More than 225,000 people are dead due to covid-19"
[…]
"step_num": 2
"action": "generate_query",
"result": {
"advanced_query": ["COVID-19 death toll October 2020", "How many people died from COVID-19 by October 2020?"]}
"step_num": 3
"action": "PureSearch",
"result": {
[…]
"step_num": 4
"action": "factuality_check",
"reasoning": "The claim states that more than 225,000 people are dead due to COVID-19 as of October 30, 2020. Evidence from the COVID Tracking Project indicates that the confirmed and probable deaths were 219,660 on October 28, 2020, which is below 225,000. Therefore, the claim is not accurate as of the specified date.",
"error": "The claim overstates the number of COVID-19 deaths as of October 30, 2020.",
"correction": "Fewer than 225,000 people are dead due to COVID-19 as of October 30, 2020.",
"factuality": "False"
"step_num": 5
"action": "feedback",
"decision": true,
"tool": ["advanced","multilingual"],
"feedback": "The evidence collected indicates that the number of confirmed and probable COVID-19 deaths was 219,660 as of October 28, 2020, which is below the claim of more than 225,000. However, the fact-checking process could benefit from generating additional questions to explore the context of the death toll further, including potential discrepancies in reporting or variations in data sources. Additionally, searching in non-English sources may provide more comprehensive data. Please reformulate questions to gather more relevant evidence."
"step_num": 6
"action": "generate_query",
"result": {
"advanced_query": ["COVID-19 death toll exceeding 225,000 by October 30, 2020",],
"multilingual_query": ["Número de muertes confirmadas por covid-19 a finales de octubre de 2020", "Número total de mortes confirmadas por covid-19 até 30 de outubro de 2020"]}
"step_num": 7
"action": "PureSearch",
"description": "Directly use searched snippets as evidences",
"result": {
[…]
"step_num": 8
"action": "factuality_check",
"reasoning": "The claim states that more than 225,000 people are dead due to COVID-19 as of October 30, 2020. Evidence from the COVID Tracking Project and many news reporting from CNN and BBC in Spanish indicates that the confirmed and probable deaths were 222,625 on October 30, 2020, which is indeed above 225,000. However, earlier evidence from October 28, 2020, indicated 219,660 deaths, which was below the claim. The new evidence confirms that the death toll exceeded 225,000 by October 30, 2020, thus supporting the claim. Therefore, the claim is factual as it aligns with the new evidence provided.",
"error": "None",
"correction": "The claim is accurate as it states that more than 225,000 people are dead due to COVID-19 as of October 30, 2020.",
"factuality": "True"
Figure 7: A simplified case due to space restriction. It shows how XLE works by providing fresh evidence different from English evidence. Complete cases can be found in our supplementary materials, where we append all of the experiment results.
A.7 Example Prompts
@@ 3. Requests for resolving vague references:
Figure 9: The prompt used in Contextual Grounding.
[Instructions]
2. Only provide domain suffixes, not full URLs.
3. Include reliable sources such as:
Government and official websites (.gov, .org)
Encyclopedia websites (.wiki)
Reputable news outlets (provide their official domain names)
6. In summary, for each claim, provide:
Recommended domain suffixes for searching
Domains to exclude (marked with a minus sign)
[response format]: [’url1’, url2’, ’-url3’]
[Examples]
<CLAIM>: The Eiffel Tower was built in 1889.
<CLAIM>: COVID-19 vaccines are safe and effective.
<CLAIM>: {input}
<HISTORY>: {feedback}
<RESPONSE>:
Figure 10: The prompt used in Credible Source Selection.
[Supported Languages]
"Dutch", "Estonian", "Filipino", "Finnish", "French (Canada)",
"Swedish", "Thai", "Turkish", "Ukrainian", "Vietnamese", "Zulu"]
[Instructions]
[response format]: [’language1’, language2’]
[Examples]
<RESPONSE>: ["German"]
<RESPONSE>: ["Korean"]
<RESPONSE>: ["Portuguese (Brazil)"]
<RESPONSE>: ["Russian", "Ukrainian"]
<RESPONSE>: None
Now complete the following:
<CLAIM>: {input}
<RESPONSE>:
Figure 11: The prompt used in Cross-Lingual Expansion.
@@ 6. In summary, for each claim, provide:
Figure 12: The prompt used in Structured Query Generation.